Circular No. 7991

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2002 S8, S9, S10, S11, T2, T3 (SOHO) Further to IAUC 7986, D. Hammer reports additional comets found

on SOHO website C2 images by D. Evans (C/2002 S8), S. Hönig (S9), M. Boschat (S10), T. Scarmato (S11), X. Leprette (T2), and M. Oates (T3). All were also observed in the C3 coronagraph except for C/2002 T2, and all were Kreutz sungrazers except C/2002 S11 (Kracht group) and S/2002 T2 (Meyer group). The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2002-T75.

Comet  2002 UT  α_{2000}  δ_{2000}
C/2002 S8 Sept.21.071  11^{h}28.^{m}4  − 0^{o}13′
C/2002 S9  23.471  11 40.2  − 0 59
C/2002 S10  26.071  11 54.4  − 1 48
C/2002 S11  30.071  12 14.6  − 1 16
C/2002 T2 Oct. 4.271  12 35.8  − 2 17
C/2002 T3  8.679  12 39.5  − 6 56

COMET P/2002 T4 (NEAT) K. Lawrence, S. Pravdo, and E. Helin, Jet Propulsion Laboratory,

report the discovery on Oct. 11.22 UT of a comet (with a nuclear condensation of diameter ∼ 4′′ and a tail ∼ 20′′ long toward the south-southwest) on NEAT images taken at Palomar. M. Hicks reports that images taken by J. Young at Table Mountain on Oct. 12.3 (through cirrus clouds) show a diffuse coma and a faint 5′′ tail to the southwest. D. Balam, University of Victoria, reports that images taken by J. Clem with the 1.82-m Plaskett telescope (also on Oct. 12.3) also show the object to be cometary in appearance (3 pixels, or 3′′.3, larger than nearby stars).

2002 UT  α_{2000}  δ_{2000}  m_{1}
Oct. 11.22228  2^{h}04^{m}33.^{s}10  +18^{o}08′40′′.9  19.3

The available astrometry (including prediscovery LINEAR observations from Oct. 4 and 9), the following elliptical orbital elements, and an ephemeris appear on MPEC 2002-T76.

T = 2002 July 25.536 TT  ω = 359.^{o}570
e = 0.44916  Ω = 358.796 }2000.0
q = 2.11594 AU  i = 6.033
a = 3.84130 AU n^{o} = 0.130914 P = 7.53 years
2002 October 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green